OpenStudy (brynndelyn):

You accept a job as an environmental engineer that pays a salary of $45,000 in the first year. After the first year, your salary increases by 3.5% per year. a. Write a rule giving your salary a sub n for your nth year of employment. b. What will your salary be during your 5th year of employment? c. What is the total amount you will earn if you work for 30 years

OpenStudy (brynndelyn):

I did a and b but I need help with c

OpenStudy (dan815):

okay can you give your expression for a and b

OpenStudy (anonymous):

leave this one to dann he's a genius

OpenStudy (dan815):

^ bahaha xD no

OpenStudy (anonymous):

Total=45000(1+3.5%)^30 Total=45000(1+0.035)^30 Total=45000(1.035)^30 Total=45000x33.79 Total=1,520,663.11